how do you do these types of equations?:
5-2(y+1)=21

Answer by rahman(247)

5-2(y+1)=21
5-2*y + (-2)*1 = 21
5 - 2y - 2 = 21
5 - 2 -2y = 21
3 - 2y = 21
-2y = 21 - 3
-2y = 18
y = 18 : (-2)
y = -9